7 Student Council President Elections from Student Counal President Elections Two candidates, A and B, are running for the president of student counail. At the Based on the results of the simulation, how likely is it that the newspapers moment, candidate A thinks that each candidate has the support of 50% of the survey will fand that ether andidate A or candidate B is supported by 60 or students. The school newspaper is going to ask a random sample of 100 students about their choice in the election. more percent of the surveyed students? To get an idea of what to expect before the survey result is published, candidate A ran a computer simulation. The simulation used a virtual population evenly split between two choices A and B from which 200 random samples of size 100 were drawn. The results of the simulation are shown in the graphic below. o A. Extremely unlikely. The probability is less than 0.05 B. Unlikely. The probablity is between 0.05 and 0.2 о С. Likely. The probability is between 0.2 and 0.8. o D. Very likely. The probability is greater than 0.8 Distribution of 200 Sample Vote Counts for Candidate A 60 50 Vote Count for Candidate A The shaded dot represents 46 vo tes for candidate A 40 45 pe here to search
Answer #1

Here the sample size n = 200

here as in the given problem we will find the points which are above 60% and below 40%.

Points above 60% = 4

points below 40% = 5

so points either above 60% or below 40% = 4 + 5 = 9

so total probability that either one will get above 60% = 9/200 = 0.045

so here this probability is less than 0.05 so the answer is extremely likely. The probability is less than 0.05. Option a is correct here.